OpenStudy (anonymous):

Get medal once again :3 What is the answer to the problem below? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@hotgurl72 B? Can you do the problem step by step for me?

OpenStudy (anonymous):

6+-4+16=?

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

@hotgurl72 Oh okay, thanks again *hug*

OpenStudy (anonymous):

No prob. Any other questions?

OpenStudy (anonymous):

@hotgurl72 Only one more 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

OpenStudy (anonymous):

Well, do you want me to tell you the answer or go through it step by step with you?

OpenStudy (anonymous):

@hotgurl72 Well tell me the answer and I figure out how to do it myself :3

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Give me a sec

OpenStudy (anonymous):

It's B again, LOL.

OpenStudy (anonymous):

@hotgurl72 That's crazy how to correct answer always seems to be B :P

OpenStudy (anonymous):

LOL. Do you want me to explain?

OpenStudy (anonymous):

Once again thanks for your help, I shall be logging off now. Oh and yes please explain before I go ^.^

OpenStudy (anonymous):

OK, so -6+18=15-3. Make sense?

OpenStudy (anonymous):

Oh I see now, either I'm a idiot in math or you're a genius :3 Anyway I be logging off now, bye *hug*

OpenStudy (anonymous):

Bye! *hug* here's my email if you have anymore questions in the future emilyaryasmith@gmail.com